OpenStudy (anonymous):

Simplify (3n - 2m)^2 = ? I will give a medal! Btw, I want an explanation, as to how to solve it, not a direct answer. ^_^ P.s. Thanks a ton!

OpenStudy (anonymous):

@AravindG @baru

OpenStudy (baru):

you have to use the "distributive property of multiplication over addition" for example \[a(b+c+d)= ab+ac+ad\]

OpenStudy (baru):

here you have\[(3n-2m)(3n-2m)\]

OpenStudy (anonymous):

i only have an a and a b

OpenStudy (anonymous):

yes, its like ab x ab

OpenStudy (anonymous):

i mean a-b x a-b :P

OpenStudy (baru):

treat the first (3n -2m) like you would have treated "a"

OpenStudy (anonymous):

... ?

OpenStudy (baru):

you get (3n-2m)3n - (3n-2m)2m

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

so, 9n-6mn - 6nm-4m ?

OpenStudy (anonymous):

so, 9n - 12mn - 4m ?

OpenStudy (anonymous):

wait, its 9n^2 - 12mn - 4m^2 isnt it?

OpenStudy (baru):

very close... but you have made a couple of errors remember that \(n \times n =n^2\) and negative number multiplied by a negative number is positive

OpenStudy (anonymous):

oh, so its 9n^2 + 12mn + 4m^2 ?

OpenStudy (baru):

everythings right, but there a mistake in the signs

OpenStudy (anonymous):

??

OpenStudy (anonymous):

what do you mean?

OpenStudy (baru):

(3n-2m)3n - (3n-2m)2m 9n^2 -6mn - (6mn - 4m^2) 9n^2 -6mn -6mn +4m^2 9n^2 -12mn +4m^2

OpenStudy (anonymous):

oh, xD i see.

OpenStudy (anonymous):

ok, thanks! ^_^

OpenStudy (baru):

no prob :)